# On point spectrum of Jacobi matrices generated by iterations of quadratic polynomials

### v1

@@ Abstract.
Jacobi matrix, point spectrum, limit periodic operator, hull, Ruelle operator
@@ 1. Introduction
In an ergodic family of operators, ergodicity implies almost sure constancy of the spectrum, as well as a.c., s.c., p.p. spectrum [17]. Historically, the a.c. spectrum was first understood in the setting of periodic operators, and the pure point spectrum gained prominence in the context of Anderson localization. Singular continuous spectrum was originally perceived as an anomaly, but then discovered to be generic in many settings [20, 3, 8]. The study of spectral type of ergodic operators is now a vast area of research; it allows only a few general results, such as Kotani theory [15, 19], and a lot of research is centered on prominent models such as the Anderson model, the Almost Mathieu operator, and the Fibonacci Hamiltonian; the dependence of spectral type on the ergodic space and the size and arithmetic properties of the parameters are of interest, and the analyses are highly model-dependent.
@@
with irrational $\omega\in{\mathbb{T}}$ is an almost periodic operator, with hull indexed by $\theta\in{\mathbb{T}}$. In the subcritical case $\lambda<1$, it has a.c. spectrum, but in the supercritical case $\lambda>1$, its spectrum is s.c. or p.p. depending on the phase [14, 12]. A remarkable result of Jitomirskaya [13] proves that the critical ($\lambda=1$) Almost Mathieu operator has empty p.p. spectrum for every choice of $\theta$, after partial results [2].
Jacobi matrices generated by iterations of quadratic polynomials are the leading model in inverse spectral theory. Starting from the pioneering work of Bellisard–Bessis–Moussa [6], they have attracted much attention, providing almost periodic operators with singular continuous spectrum; cf. [4, 5, 7, 18]. In this paper, we prove absence of pure point spectrum in the hull for this model and view this as a counterpart of Jitomirskaya’s theorem.
To describe the model, start with an even quadratic polynomial $T(z)=z^{2}-\lambda$, $\lambda>2$. Its Julia set
@@
uniformly in $k$ and $s$. Due to this property the sequence can be extended to a continuous function on dyadic integers ${\mathbb{Z}}_{2}$
@@
This extends the half-line Jacobi matrix associated to $\mu_{\mathsf{E}}$ to a full-line limit-periodic Jacobi matrix $J_{0}$, and describes the hull of this matrix as $\{J_{\varkappa}\mid{\varkappa}\in{\mathbb{Z}}_{2}\}$ where
@@
The key tool are matrix Ruelle operators, see Definition 2.3. A matrix generalization of Ruelle theory [9] seems a hard problem, due to noncommutativity. For instance, whereas the multiplicative representation of Schur functions, after separation of a Blaschke product, is a simple observation, its matrix valued analog is the multiplicative theory of $J$-contractive matrices [1]. Note that de Branges theory of canonical systems [10] is its partial case. However, we were able to reduce our problem to matrix Ruelle operators with special rational weights, and due to this, our construction becomes parallel to that suggested in [21].
It is an open problem whether the spectral measure corresponding to an element $\varkappa$ with periodic digits, such as $\varkappa=1+\sum_{j=0}^{\infty}2^{2j}$ (in ${\mathbb{Z}}_{2}$), is mutually singular to the spectral measure of $J_{0}$. The operator $J_{\varkappa}$ corresponds to a matrix Ruelle operator, and using Ruelle theory for analogy suggests that different weights should correspond to mutually singular measures. If so, this would be another example of operators in the same hull which are not unitarily equivalent, in strong contrast to the results of [22] for reflectionless operators on regular Widom spectra with DCT property, where all operators in the same hull are unitarily equivalent to each other.
@@
Some explicit formulas are omitted from the journal version of this paper, and are included in the arXiv version.
We thank David Damanik for helpful information.
@@ 2.1. Background: Ruelle operators and half-line Jacobi matrices
with a positive continuous function $h\in C({\mathsf{E}})$ and a (positive) measure $\nu$. In particular, the equilibrium measure $\mu_{\mathsf{E}}$ for the set ${\mathsf{E}}$ is an eigenvector for ${\mathcal{L}}_{0}^{*}$,
@@
Since $a_{0}=0$, this gives a direct sum decomposition $J_{0}=J_{+}\oplus J_{-}$. We use +=${\mathbb{Z}}_{+}=\{n\in{\mathbb{Z}}\mid n\geq 0\}$ and ${\mathbb{Z}}_{-}={\mathbb{Z}}\setminus{\mathbb{Z}}_{+}$.
It was proved in [21] that the canonical spectral measure $\nu(dx)$ of $J_{-}$ is the eigenmeasure of the adjoint to the Ruelle operator
@@
i.e., ${\mathcal{L}}_{2}^{*}\nu=\rho_{2}\nu$ and $\rho_{2}=\frac{1}{2a_{-1}^{2}}$. Related orthogonal polynomials $P_{n}(x,\nu)$ possess the renorm relation
@@
2.2. Shift in ${\mathbb{Z}}_{2}$ and associated to $J_{\varkappa}$ matrix Ruelle operators
@@
Let $Ve_{n}=e_{2n}$. The following identity holds
@@
Due to (2.1) the operator $f(x)\mapsto f(T(x))$ is an isometry in $L^{2}_{\mu_{{\mathsf{E}}}}$. Its adjoint is $\frac{1}{2}{\mathcal{L}}_{0}$ viewed as an operator in $L^{2}_{\mu_{{\mathsf{E}}}}$. With the standard eigenfunction expansion $\ell^{2}({\mathbb{Z}}_{+})\to L^{2}_{\mu_{\mathsf{E}}}$, $e_{n}\mapsto p_{n}(\cdot,\mu)$, the operator $f\mapsto f\circ T$ is conjugated to the operator $V$ on $\ell^{2}({\mathbb{Z}}_{+})$. Thus the identity in the spectral representation
$\frac{1}{2}\sum_{T(y)=x}\frac{f(T(y))}{y-z}=\frac{1}{2}\frac{T^{\prime}(z)}{x-T(z)}f(x)$
@@
$V^{*}(J_{+}-z)^{-1}V=\frac{1}{2}T^{\prime}(z)(J_{+}-T(z))^{-1}.$
Likewise, the standard eigenfunction expansion for $J_{-}$ can be written as $\ell^{2}({\mathbb{Z}}_{-})\to L^{2}_{\nu}$, $e_{-1-n}\mapsto p_{n}(\cdot,\nu)$, and due to (2.2), the adjoint of $V$ on $\ell^{2}({\mathbb{Z}}_{-})$ in the spectral representation takes the form
$a_{-1}{\mathcal{L}}_{1}f(x)=a_{-1}\sum_{T(y)=x}\frac{f(y)}{T^{\prime}(y)},\quad f\in L^{2}_{\nu}.$
Similarly to $J_{+}$, the identity
@@
$V^{*}(J_{0}-z)^{-1}V=\frac{1}{2}T^{\prime}(z)(J_{0}-T(z))^{-1}$
for operators on $\ell^{2}({\mathbb{Z}})$, and since $VS=S^{2}V$ we have
@@
The following renorm equation holds
@@
where $p_{k}^{\pm},q_{k}^{\pm}$ are orthogonal polynomials of the first and second type generated by $(J_{\varkappa})_{\pm}$. In particular
$e_{-2}\mapsto\begin{pmatrix}\frac{x}{a_{{\varkappa}-1}}\\ -\frac{a_{{\varkappa}}}{a_{{\varkappa}-1}}\end{pmatrix}=\begin{pmatrix}\frac{T^{\prime}(x)}{2a_{{\varkappa}-1}}\\ -\frac{a_{{\varkappa}}}{a_{{\varkappa}-1}}\end{pmatrix},\quad e_{0}\mapsto\begin{pmatrix}0\\ 1\end{pmatrix}.$
@@
where $g$ is $2\times 2$ continuous matrix function on ${\mathsf{E}}$. Based on this we define matrix Ruelle operators.
@@
and $g(x)\geq 0$ implies $({\mathcal{M}}_{\varkappa}g)(x)\geq 0$.
2.3. Action of Ruelle operators on a class of rational functions
In this subsection we show that the collection of rational functions with simple poles at $T^{\circ k}(0)$ is invariant for the action of ${\mathcal{M}}_{\varkappa}$ (at the moment we do not consider its closure).

### v2

@@ Abstract.
Jacobi matrix, point spectrum, limit periodic operator, hull, Julia set, Ruelle operator
@@ 1. Introduction
In an ergodic family of operators, ergodicity implies almost sure constancy of the spectrum, as well as a.c., s.c., p.p. spectrum [20]. Historically, the a.c. spectrum was first understood in the setting of periodic operators, and the pure point spectrum gained prominence in the context of Anderson localization. Singular continuous spectrum was originally perceived as an anomaly, but then discovered to be generic in many settings [23, 3, 8]. The study of spectral type of ergodic operators is now a vast area of research; it allows only a few general results, such as Kotani theory [18, 22], and a lot of research is centered on prominent models such as the Anderson model [12], the Almost Mathieu operator [16], and the Fibonacci Hamiltonian [10]; the dependence of spectral type on the ergodic space and the size and arithmetic properties of the parameters are of interest, and the analyses are highly model-dependent.
@@
with irrational $\omega\in{\mathbb{T}}$ is an almost periodic operator, with hull indexed by $\theta\in{\mathbb{T}}$. In the subcritical case $\lambda<1$, it has a.c. spectrum, but in the supercritical case $\lambda>1$, its spectrum is s.c. or p.p. depending on the phase [17, 14]. A remarkable result of Jitomirskaya [15] proves that in the critical case $\lambda=1$, Almost Mathieu operator has empty p.p. spectrum for every choice of $\theta$, after partial results [2].
Jacobi matrices generated by iterations of quadratic polynomials are the leading model in inverse spectral theory. They have attracted much attention, providing almost periodic operators with zero Lebesgue measure spectrum; cf. [6, 4, 5, 7, 21]. In this paper, we prove absence of pure point spectrum in the hull for this model and view this as a counterpart of Jitomirskaya’s theorem.
To describe the model, start with an even quadratic polynomial
$T(z)=z^{2}-\lambda,\qquad\lambda>2.$
Its Julia set
@@
uniformly in $k$ and $s$. Due to this property the sequence can be extended to a continuous function on the space of dyadic integers ${\mathbb{Z}}_{2}$
@@
(recall that ${\mathbb{Z}}_{2}$ is the completion of $\mathbb{Z}$} with respect to the metric $d(m,n)=1/2^{j}$ if $m-n\in 2^{j}{\mathbb{Z}}\setminus 2^{j+1}{\mathbb{Z}}$). This extends the half-line Jacobi matrix associated to $\mu_{\mathsf{E}}$ to a full-line limit-periodic Jacobi matrix $J_{0}$, and describes the hull of this matrix as $\{J_{\varkappa}\mid{\varkappa}\in{\mathbb{Z}}_{2}\}$ where
@@
The key tool are matrix Ruelle operators, see Definition 2.7. A matrix generalization of Ruelle theory [9] seems to be a hard problem, due to noncommutativity. For instance, whereas the multiplicative representation of Schur functions, after separation of a Blaschke product, is a simple observation, its matrix valued analog is the multiplicative theory of $J$-contractive matrices [1], which includes the de Branges theory of canonical systems [11]. However, we were able to reduce our problem to matrix Ruelle operators with special rational weights. This approach motivates a more systematic study of matrix Ruelle operators.
Dyadic integers have series expansions ${\varkappa}=\sum_{k=0}^{\infty}{\epsilon}_{k}2^{k}$, ${\epsilon}_{k}\in\{0,1\}$, and in our analysis, it is necessary to separate the hull into elements for which the dyadic expansion contains stretches of zeros and ones of uniformly bounded length, ${\mathcal{F}}_{\sigma}$, and its complement, ${\mathcal{G}}_{\delta}$. Different arguments are used in the two cases.
It is an open problem whether the spectral measure corresponding to an element $\varkappa$ with periodic digits, such as $\varkappa=1+\sum_{k=0}^{\infty}2^{2k+1}$ (in ${\mathbb{Z}}_{2}$), is mutually singular to the spectral measure of $J_{0}$. The operator $J_{\varkappa}$ corresponds to a matrix Ruelle operator, and using Ruelle theory for analogy suggests that different weights should correspond to mutually singular measures. If so, this would be another example of operators in the same hull which are not unitarily equivalent, in strong contrast to the results of [25] for reflectionless operators on regular Widom spectra with DCT property, where all operators in the same hull are unitarily equivalent to each other.
@@
Structure of the paper: In Section 2 we introduce matrix Ruelle operators associated to the hull $\{J_{\varkappa}\}$. We divide the hull into elements for which $\varkappa$} contains stretches of zeros and ones of uniformly bounded length, ${\mathcal{F}}_{\sigma}$, and its complement, ${\mathcal{G}}_{\delta}$. In Section 3 we prove Theorem 1.1 in the generic case, i.e., ${\varkappa}\in{\mathcal{G}}_{\delta}$. The case ${\varkappa}\in{\mathcal{F}}_{\sigma}$ requires uniform boundedness from below of a certain sequence of matrix valued test functions, which is proved in Section 4. With these bounds we conclude the proof of Theorem 1.1 in Section 4.
Acknowledgment: We thank David Damanik for helpful information. P.Y. thanks Paul F.X. Müller for useful discussions.
@@ 2.1. Background: Ruelle operators and half-line Jacobi matrices
with a positive continuous function $h\in C({\mathsf{E}})$ and a (positive) measure $\nu$. With a slight abuse of notation, we will denote by ${\mathcal{L}}_{j}$, $j=0,1,2$ the Ruelle operators
$({\mathcal{L}}_{j}f)(x)=\sum_{T(y)=x}\frac{f(y)}{T^{\prime}(y)^{j}}.$
In particular, the equilibrium measure $\mu_{\mathsf{E}}$ for the set $\mathsf{E}$} is an eigenvector for ${\mathcal{L}}_{0}^{*}$,
@@
Since $a_{0}=0$, this gives a direct sum decomposition $J_{0}=J_{+}\oplus J_{-}$. We use ${\mathbb{Z}}_{+}=\{n\in{\mathbb{Z}}\mid n\geq 0\}$ and $\mathbb{Z}$$_{-}$$\mathbb{Z}$$\setminus{\mathbb{Z}}_{+}$. Note that all diagonal Jacobi coefficients are zero.
It was proved in [24] that the canonical spectral measure $\nu(dx)$ of $J_{-}$ is the eigenmeasure of the adjoint of the Ruelle operator
@@
i.e., ${\mathcal{L}}_{2}^{*}\nu=\rho_{2}\nu$ and $\rho_{2}=\frac{1}{2a_{-1}^{2}}$. The corresponding orthogonal polynomials $p_{n}(x,\nu)$ possess the renormalization relation
@@
2.2. Shift in ${\mathbb{Z}}_{2}$ and matrix Ruelle operators associated to $J_{\varkappa}$
We will use the operator $V$ on $\ell^{2}({\mathbb{Z}})$ given by $Ve_{n}=e_{2n}$, and denote by $V_{\pm}$ its restrictions to $\ell^{2}({\mathbb{Z}}_{\pm})$.
@@ 2.2. Shift in ${\mathbb{Z}}_{2}$ and matrix Ruelle operators associated to $J_{\varkappa}$
For all $z\in{\mathbb{C}}\setminus{\mathsf{E}}$,
@@
Due to (2.1) the operator $f(x)\mapsto f(T(x))$ is an isometry in $L^{2}_{\mu_{{\mathsf{E}}}}$. Its adjoint is $\frac{1}{2}{\mathcal{L}}_{0}$ viewed as an operator in $L^{2}_{\mu_{{\mathsf{E}}}}$. With the standard eigenfunction expansion $\ell^{2}({\mathbb{Z}}_{+})\to L^{2}_{\mu_{\mathsf{E}}}$, $e_{n}\mapsto p_{n}(\cdot,\mu)$, the operator $f\mapsto f\circ T$ is conjugated to $V_{+}$. Thus the identity in the spectral representation
$\frac{1}{2}\sum_{T(y)=x}\frac{f(T(y))}{y-z}=\frac{1}{2}\frac{T^{\prime}(z)}{x-T(z)}f(x)$ (2.4)
@@
$V_{+}^{*}(J_{+}-z)^{-1}V_{+}=\frac{1}{2}T^{\prime}(z)(J_{+}-T(z))^{-1}.$
Likewise, the standard eigenfunction expansion for $J_{-}$ can be written as $\ell^{2}({\mathbb{Z}}_{-})\to L^{2}_{\nu}$, $e_{-1-n}\mapsto p_{n}(\cdot,\nu)$, and due to (2.2), the adjoint of $V_{-}$ in the spectral representation takes the form
$a_{-1}{\mathcal{L}}_{1}f(x)=a_{-1}\sum_{T(y)=x}\frac{f(y)}{T^{\prime}(y)},$
used on continuous functions $f$ and extended by continuity to $L^{2}_{\nu}$. Similarly to $J_{+}$, the identity
@@
$V^{*}(J_{0}-z)^{-1}V=\frac{1}{2}T^{\prime}(z)(J_{0}-T(z))^{-1}.$
Using the shift operator $Se_{n}=e_{n+1}$ and noting $VS=S^{2}V$, we have
@@
We also introduce a modified shift on ${\mathbb{Z}}_{2}$ defined by
which naturally appears in the following relation:
@@
The following renormalization equation holds
@@
where $p_{k}^{\pm},q_{k}^{\pm}$ are orthogonal polynomials of the first and second kind generated by the half-line restrictions $(J_{\varkappa})_{\pm}$ of $J_{\varkappa}$ to $\ell^{2}({\mathbb{Z}}_{\pm})$. In particular
$e_{-2}\mapsto\begin{pmatrix}\frac{x}{a_{{\varkappa}-1}}\\ -\frac{a_{{\varkappa}}}{a_{{\varkappa}-1}}\end{pmatrix}=\begin{pmatrix}\frac{T^{\prime}(x)}{2a_{{\varkappa}-1}}\\ -\frac{a_{{\varkappa}}}{a_{{\varkappa}-1}}\end{pmatrix},\quad e_{0}\mapsto\begin{pmatrix}0\\ 1\end{pmatrix}$. (2.6)
@@
Remark 2.3.
The modified shift obeys
${\varkappa}=\begin{cases}2\hat{\underline{s}}{\varkappa}&{\varkappa}\text{ even}\\ -1+2\hat{\underline{s}}{\varkappa}&{\varkappa}\text{ odd}\end{cases}$
so this is equivalent to the standard shift applied to $\varkappa$}. In other words
$\hat{\underline{s}}{\varkappa}=-\underline{s}(-{\varkappa}).$ (2.10)
Definition 2.4.
We denote
$\hat{\underline{s}}^{n}{\varkappa}=\sum_{k=0}^{\infty}{\epsilon}^{n}_{k}2^{k}.$
Define
$\kappa=\kappa({\varkappa})=\sum_{n=0}^{\infty}{\epsilon}^{n}_{0}2^{n}.$
It gives the identity related to the standard shift
$\kappa(\hat{\underline{s}}{\varkappa})=\underline{s}\kappa({\varkappa}).$
We decompose ${\mathbb{Z}}_{2}$ into two shift invariant subsets.
Definition 2.5.
For an arbitrary $N\in{\mathbb{Z}}_{+}$ we define ${\mathcal{F}}_{N}$ as a collection of ${\varkappa}\in{\mathbb{Z}}_{2}$ such that $\varkappa$} has at most $N$ consecutive zeros or ones in its dyadic representation.
Lemma 2.6.
Let ${\varkappa}\in{\mathcal{F}}_{N}$. Then $\kappa({\varkappa})\in{\mathcal{F}}_{N+1}$.
Proof.
We will prove an explicit relation (2.11) between components of $\varkappa$} and $\kappa({\varkappa})$. Recall that
$\hat{\underline{s}}^{n}{\varkappa}=-\underline{s}^{n}(-{\varkappa}).$
Further, let ${\varkappa}=\sum_{k\geq 0}{\epsilon}_{k}2^{k}$. Since
$\sum_{k\geq 0}{\epsilon}_{k}2^{k}+\sum_{k\geq 0}(1-{\epsilon}_{k})2^{k}=-1$
we have
${\varkappa}=1+\sum_{k\geq 0}(1-{\epsilon}_{k})2^{k}.$
Let ${\varkappa}\in{\mathcal{F}}_{N}$. Let $n_{0}=\min\{k\in{\mathbb{Z}}_{+}\mid{\epsilon}_{k}=1\}$. Note $n_{0}\leq N$. By repeatedly applying the modified shift, we see that ${\epsilon}_{0}^{n}=0$ for all $n<n_{0}$ and ${\epsilon}_{0}^{n_{0}}=1$. Applying $n_{0}$ shifts we get an odd number
$\hat{\underline{s}}^{n_{0}}{\varkappa}=1+\sum_{k\geq 1}{\epsilon}_{k+n_{0}}2^{k},\quad n_{0}\leq N.$
Then
$\hat{\underline{s}}^{n_{0}}{\varkappa}=1+\sum_{k\geq 1}(1-{\epsilon}_{k+n_{0}})2^{k}$
and
$\underline{s}^{n}(-\hat{\underline{s}}^{n_{0}}{\varkappa})=\sum_{k\geq 0}(1-{\epsilon}_{k+n_{0}+n})2^{k},\quad n\geq 1.$
We obtain
$\hat{\underline{s}}^{n+n_{0}}{\varkappa}=1+\sum_{k\geq 0}{\epsilon}_{k+n+n_{0}}2^{k},$
from which we read off all the remaining digits of $\kappa({\varkappa})$ to be
${\epsilon}^{n+n_{0}}_{0}={\epsilon}_{n+n_{0}}+1\mod 2,\quad n\geq 1.$ (2.11)
Note that ${\epsilon}^{n_{0}}_{0}=1$ and ${\epsilon}^{n}_{0}=0$ for $n<n_{0}$. So, if ${\epsilon}_{k+n_{0}}=0$ for $1\leq k\leq N$ we get $N+1$ consecutive ones in the decomposition of $\kappa({\varkappa})$. ∎
Evidently, each ${\mathcal{F}}_{N}$ is a perfect closed set. We denote
${\mathcal{F}}_{\sigma}=\cup_{N\geq 1}{\mathcal{F}}_{N},\quad{\mathcal{G}}_{\delta}={\mathbb{Z}}_{2}\setminus{\mathcal{F}}_{\sigma}.$
@@
where $g$ is $2\times 2$ continuous matrix function on $\mathsf{E}$}. Motivated by this duality and Lemma 2.2, we define matrix Ruelle operators.
@@
and $g(x)\geq 0$ implies $({\mathcal{M}}_{\varkappa}g)(x)\geq 0$. In particular (2.5) has the form
$\Sigma_{{\varkappa}}(dx)=({\mathcal{M}}_{{\varkappa}}^{*}\Sigma_{\hat{}\underline{s}{\varkappa}})(dx).$ (2.13)
3. Main theorem: generic case
Recall that
${\mathsf{E}}\subset[-\xi,\xi],\quad\xi=\frac{1+\sqrt{1+4\lambda}}{2},$
and $\pm\xi\in{\mathsf{E}}$. In fact, $T^{-1}([-\xi,\xi])\subset[-\xi,\xi]$ and ${\mathsf{E}}=\cap_{n=0}^{\infty}(T^{-1})^{\circ n}([-\xi,\xi])$. Moreover, note that $T(0)=-\lambda<-\xi$.
First we show that the collection of rational functions with simple poles at $T^{\circ k}(0)\in{\mathbb{R}}\setminus[-\xi,\xi]$ is invariant for the action of ${\mathcal{M}}_{\varkappa}$ (we do not consider its closure).